OpenStudy (moongazer):

Max.\:\:Height = \frac{-(Vi \sin\theta)^2}{2g}

OpenStudy (moongazer):

is\[Max.\:\:Height = \frac{-(Vi \sin\theta)^2}{2g}\] the same as \[Max.\:\:Height = \frac{u^2\sin^2\theta}{2g}\]

OpenStudy (moongazer):

???

OpenStudy (anonymous):

what does i stands for

OpenStudy (moongazer):

initial velocity

OpenStudy (moongazer):

also the u in 2nd fomula is initial velocity

OpenStudy (anonymous):

same

OpenStudy (moongazer):

because that is the formula my teacher gave me

OpenStudy (anonymous):

it is correct

OpenStudy (anonymous):

and same as the one I gave you

OpenStudy (anonymous):

I use u for initial velocity and v for final

OpenStudy (moongazer):

ok thanks

OpenStudy (anonymous):

moongrazer how did your exam go

OpenStudy (anonymous):

for university

OpenStudy (moongazer):

w8

OpenStudy (moongazer):

why is it i am having a different answer?

OpenStudy (moongazer):

could you check it?

OpenStudy (moongazer):

i am answering the question here http://openstudy.com/groups/physics#/groups/physics/updates/4e4cc3070b8b9588049ae330

OpenStudy (anonymous):

I will try

OpenStudy (moongazer):

when i use the formula my teacher gave me it is 52.70 and if i tried yours it is 43.63

OpenStudy (anonymous):

I am getting 51.7

OpenStudy (anonymous):

try it again must be missing something it is square

OpenStudy (moongazer):

on both?

OpenStudy (anonymous):

no max height I am talking about max height let me do range too

OpenStudy (moongazer):

no, i mean do you get 51.7 on using both formulas?

OpenStudy (anonymous):

246.20 range 51.7 max height

OpenStudy (moongazer):

did you used the formula my teacher gave me?

OpenStudy (anonymous):

it's the same answer is same 51.7 and 246.20 I am using calculator the angle is 40 and initial velocity is 50

OpenStudy (anonymous):

^^^^^^^^^even with your teacher's formula

OpenStudy (moongazer):

ohhh i had the correct answer now

OpenStudy (anonymous):

good job : )

OpenStudy (moongazer):

i squared 50 that is why i got a wrong answer

OpenStudy (moongazer):

Thanks for helping me :)

OpenStudy (anonymous):

anytime : )

OpenStudy (moongazer):

physics have a lot of formulas

OpenStudy (moongazer):

We have an index card that contains i think more or less 10 formulas that we will use for our exam. But i think we will not gonna use it all. ^_^

OpenStudy (anonymous):

Good Luck ...just practice questions and then you won't have problem remembering formulas

OpenStudy (moongazer):

yup, practices makes perfect

OpenStudy (moongazer):

i have a last question

OpenStudy (anonymous):

okay I will try that in 5 minutes I'm hungry now gotta eat somthing

OpenStudy (moongazer):

how do you transform an equation if you want to have the formula for\[\theta\] like \[R= (-Vi^2 \sin \theta)/g\] please explain how thanks for helping me :) \[Vi = initial velocity\]

OpenStudy (moongazer):

is it the same as\[R = v^2 \sin (2\theta)/g\]

OpenStudy (moongazer):

??

OpenStudy (anonymous):

Okay I am back. \[R = \frac{V_i^2 sin2\theta}{g}\] \[R \times g = V_i^2 sin2\theta\] \[\frac{R \times g }{V_i^2}=sin2 \theta\] \[sin^{-1} \frac{R \times g }{V_i^2} = 2\theta\] \[\frac{1}{2}sin^{-1} \frac{R \times g }{V_i^2} = \theta\]